if f(x)=2x^2+5 sqrt(x-2) complete the following statement(round your answer to the nearest hundredth): f(0)=
Answer by rothauserc(4718)

f(x) = 2x^2 +5sqrt(x-2)
:
f(0) = 2(0)^2 +5sqrt(0-2)
:
f(0) = 5sqrt(-2)
:
f(0) = 5 * i * sqrt(2)
:
Note i = sqrt(-1)
:
f(0) = 5 * i * 1.4142
:
f(0) = 7.071 * i
:
************
f(0) = 7.07i
************
: